OpenStudy (anonymous):

f(x) = 7x(3x +8) using the product rule for derivatives. please help :)

OpenStudy (amistre64):

7(3x+8) + 7x(3)

OpenStudy (amistre64):

21x +56 +7x + 21 28x + 77

OpenStudy (amistre64):

7(4x +11)

OpenStudy (amistre64):

uzma...be nice and play along :)

OpenStudy (anonymous):

mean?:)

OpenStudy (anonymous):

i gave u a chance to complete :P

OpenStudy (amistre64):

lol

OpenStudy (anonymous):

was checking ur computations;)

OpenStudy (amistre64):

my computations are full of errors :)

OpenStudy (anonymous):

yup :P

OpenStudy (anonymous):

so should i start?

OpenStudy (amistre64):

be my guest :)

OpenStudy (anonymous):

oh so nice of u :)

OpenStudy (anonymous):

do you all do this all day? :)

OpenStudy (amistre64):

point put that 7x(3) does not equal 7x + 21 lol

OpenStudy (anonymous):

yea i got it

OpenStudy (anonymous):

parison ...whenever i get free time, its a healthy activity :)

OpenStudy (amistre64):

half the day; the other half is spent foraging for math books and laundering vitamin pills

OpenStudy (anonymous):

definitely. i am taking an online statistics course & am so happy I found this site

OpenStudy (anonymous):

:D yea

OpenStudy (amistre64):

42x+56....

OpenStudy (amistre64):

7(6x+8)

OpenStudy (anonymous):

so f'(x)= (7x)'(3x+8)= 7x(3x+8)'

OpenStudy (anonymous):

where ' denotes the derivative

OpenStudy (amistre64):

= means + when unless it means = lol

OpenStudy (anonymous):

f'(x)=7(3x+8)+7x(3)=21x+56+21x=42x+56=7(6x+8)